CHAPTER 8 QUESTION

Use the data provided to answer the questions that follow relating to Newtech Ltd for 2017:

Statement of Financial Position:

Assets                                                         2017                                            2016

Non-Current/Fixed                                  4 200 000                                   3 000 000

Inventory                                                     400 000                                      600 000

Receivables                                              1 550 000                                  1 200 000

Cash                                                             600 000                                     300 000

                                                                   6 750 000                                 5 100 000

Equity and Liabilities

Share Capital (R2 shares)                           4 200 000                                4 000 000

Retained Income                                            600 000                                   300 000

Long term Debt                                              250 000                                    200 000

Payables                                                      1 700 000                                   600 000

                                                                     6 750 000                                5 100 000

The Statement of Comprehensive Income for the year ended 2017:

Sales (10% on credit)                                  10 200 000

Cost of sales (80% on credit)                        4 080 000

Expenses                                                      3 200 000

Net Income after Tax                                     2 000 000

Dividends                                                      1 700 000

Retained Income                                          3 000 000

NB: Shares are currently trading at R2, 80 per share.

Required:

1.1. Calculate the net profit on turnover for 2015. (2)

1.2. Calculate the current ratio for both years. (4)

1.3. Calculate the acid test ratio for both years. (4)

1.4. Accounts Payable Turnover

1.4.1. Calculate the accounts payable period (in days), noting that Newtech Ltd has, after tough negotiations secured a 90 day account with all its creditors.

Note: Use average accounts payable for this calculation. (4)

1.4.2. Offer constructive advice to Newtech Traders as to whether this is an ideal situation. (3)

1.5. Calculate the return on equity. Explain if the shareholders will be satisfied with the return. (4)

1.6. Calculate the inventory turnover ratio (use average inventory) and explain the significance of this ratio.
Answer #1
2017 2016
1.1) Net profit on turnover = 2000000/10200000 = 19.61%
1.2) Current Ratio:
= (400000+1550000+600000)/1700000 = 1.50
= (600000+1200000+300000)/600000 = 3.50
1.3) Acid test ratio:
= (1550000+600000)/1700000 = 1.26
= (1200000+300000)/600000 = 2.50
1.4) Accounts payable turnover:
= 4080000/((1700000+600000)/2)) = 3.55
1.4.1) Accounts payable period = 365/3.55 = 103 Days
1.4.2) As against the negotiated period the actual payment period is higher at 103 days
which is not an ideal situation. Delaying payments past due will affect relationship
with suppliers.
1.5) ROE = 2000000*((4800000+4300000)/2)) = 43.96%
Yes, the shareholders should be satisfied with the return.
1.6) Inventory turnover = 4080000/((400000+600000)/2)) = 8.16
The ratio tell us how many times inventory is turned over by cost
of goods sold. More the turnover more the efficiency.